# A comprehensive analysis toward the Fermi-LAT source 4FGL J1846.90227 Jets of a proto-planetary nebula producing -rays?

### v1

@@ 4.4 What about MYSO MSX G30?
From our analysis based on ALMA data, we found that the confirmed MYSO MSX G30 is associated with a dense molecular core, which presents evidence of star formation activity due to the presence of numerous molecular lines (Fig.ā4), which are characteristic of hot molecular cores. This chemical richness is an interesting issue to be studied in deep in a future work. In particular, what is most relevant to this study is the detection of formaldehyde molecular outflows (see Fig.ā5), which could explain the -ray emission of the Fermi-LAT source, as it was proposed by Munar-Adrover etĀ al. (2011). However, based on the position of MYSO MSX G30 at the border of the confidence ellipse of the Fermi-LAT source, and the fact that the MYSO does not present any X-ray or centimeter radio emissions associated, we suggest that, it is very likely that the molecular outflows detected towards this MYSO are not energetic enough to be responsible for the gamma-ray emission. Therefore, we conclude that it is very unlikely that the counterpart of the high-energy source is such a young protostar.
4.5 What about IRAS 18443-0231?
@@
The non-thermal nature of the radio continuum emission from IRAS 18443-0231 and the morphology of the continuum emission at 3 GHz indicate the presence of jets and accelerated particles. Additionally, we discovered a red-shifted molecular outflow supporting the presence of such jets in the region, which is in agreement with the detected H2O maser (Urquhart etĀ al., 2011).
@@
Taking into account that this source is cataloged as a PN candidate, our findings indicate that IRAS 18443-0231 actually would be a proto-planeatary nebula (pPN), a short-lived transition (about yr) from the asymptotic giant branch (AGB) to the PN phase. There are not much cases that non-thermal radio continuum emission was observed toward this kind of source (see the case of IRAS 15103-5754; SuĆ”rez etĀ al. 2015), and on the other side, it is more common observing molecular outflows in pPNe (Cox etĀ al., 2000; Lorenzo etĀ al., 2021). The non-thermal radio emission in pPNe suggests that the emitting electrons arise at collisions between the fast and slow AGB winds that are observed predominantly on the front sides of the circumstellar shells (Bains etĀ al., 2009).
@@
Having discarded all the sources presented above as responsible of the gamma-ray emission, taking into account the location of IRAS 18443-0231 (almost at the center of the Fermi confidence ellipse), and mainly considering the presence of non-thermal jets accelerating particles, we wonder if IRAS 18443-0231 could be the counterpart of 4FGL J1846.90227.
@@
As mention by SuĆ”rez etĀ al. (2015) for IRAS 15103-5754, the radio spectral index 0.08 measured in IRAS 18443-0231 is similar to that found in synchrotron emission in galaxies with active nuclei. Following the authors, in the energy conditions of PNe, mass-loss processes can produce collisionless, non-relativistic shocks, and electrons may undergo diffusive shock acceleration up to very high velocities (āfirst-order Fermi accelerationā). In such a scenario, in which particles are accelerated, and in addition the source seems to be embedded in dense molecular gas (see Sect.ā3.5), processes such as proton-proton collisions and relativistic Bremsstrahlung are likely to occur generating gamma-rays. In that case, IRAS 18443-0231 would be the responsible of the gamma-ray emission. If this is the case, it would be the first reported pPN related to very high energy emission, and hence, multiwavelength dedicated observations and modeling are necessary to understand the mechanisms of gamma-rays production.
@@ 5 Summary and concluding remark
In previous works the nature of the Fermi-LAT source 4FGL J1846.90227 was discussed. In one of them it was suggested to be a blazar, and in other one it was mentioned the presence of a massive protostar within the Fermi confidence ellipse suggesting a possible association. Considering this discrepancy, in this work, we analyzed several counterpart candidates to the gamma-ray source based on a multiwavelength analysis.
@@
First of all, from JVLA data we determined the thermal nature of a new extended radio source in the field, which indicates that this uncatalogued radio continuum source is an HiiĀ region and not a supernova remnant, ruling out that such a source is responsible for the gamma-ray emission. Using ALMA data, we confirmed that MSX6C G030.198100.1691 is a massive protostar with molecular outflow activity. However, since its peripheral location respect to the Fermi 95% confidence ellipse, and the lack of associated X-ray and centimeter radio continuum emissions, it is very unlikely that this massive protostar is responsible of the gamma-ray emission.
@@
We found that sources SwXF4 J184651.6-022507 and SwXF4 J184650.7-022904, previously suggested to be blazars, do not meet the mid-IR color-color criteria for this kind of sources. Additionally, they lie almost at the galactic plane, which weakens the possibility of a blazar detection. Moreover, the second one is related to the 4XMM 184650.6022907 source, and based on XMM-Newton data, we found that this X-ray source, which lies closer to the center of the Fermi 95% confidence ellipse, is clearly associated with the spectroscopic binary star BD-02 4739. The X-ray spectrum of 4XMM 184650.6022907 was well fitted by an optically thin thermal plasma model with a relatively low X-ray luminosity of about 1030 erg s-1. The soft thermal nature of the X-ray spectrum discards that 4XMM 184650.6022907 is associated with a blazar. In fact, the X-ray spectrum is consistent with a white-dwarf (WD) symbiotic system. Thus, taking into account the results of the analysis of the X-ray data and the optical/infrared SED, we conclude that BD02 4739 is a non-burning shell -type WD symbiotic system with a sub-giant donor star. The way that this source can be considered as the responsible for the gamma emission is that it had experimented a nova event. Our findings discard such a possibility.
@@
Finally, we found that IRAS 18443-0231, cataloged as a PN candidate located at a distance of about 8 kpc, is actually a protoplanetary nebula. IRAS 18443-0231 is the closest source in the region to the center of the Fermi 95% confidence ellipse. We discovered non-thermal radio continuum emission toward such a source (spectral index ), indicating syncrotron emission likely due to particles accelerated in jets. Additionally, we discovered a molecular outflow, supporting the presence of jets, and we found that the source is embedded in a dense molecular clump. IRAS 18443-0231 presents hard X-ray extended emission, with the same inclination as the molecular outflow.
@@
Considering all the presented results, we conclude that the most promising source to be associated with 4FGL J1846.90227 is IRAS 18443-0231. The detection of syncrotron emission indicate that particles are accelerated at high velocities in the region, and considering that source is embedded in a dense molecular clump, we suggest that processes such as proton-proton collisions and relativistic Bremsstrahlung are likely to occur. If this is the case, it would be the first reported protoplanetary nebula related to very high energy emission, and hence, multiwavelength dedicated observations and modeling are necessary to understand the mechanisms of gamma-rays production in this kind of source.
@@ Acknowledgments
M.O., A.P., and S.P. are members of the Carrera del Investigador CientĆfico of CONICET, Argentina. This work was partially supported by the Argentina grants PIP 2021 11220200100012 and PICT 2021-GRF-TII-00061 awarded by CONICET and ANPCYT. This work is based on the following ALMA data: ADS/JAO.ALMA 2015.1.01312. ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), MOST and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ.

### v2

@@ 4.4 What about MYSO MSX G30?
From our analysis based on ALMA data, we found that the confirmed MYSO MSX G30 is associated with a dense molecular core, which presents evidence of star formation activity due to the presence of numerous molecular lines (Fig.ā4), which are characteristic of hot molecular cores. This chemical richness is an interesting issue to be studied in deep in a future work. In particular, what is most relevant to this study is the detection of formaldehyde molecular outflows (see Fig.ā5), which could explain the -ray emission of the Fermi-LAT source, as it was proposed by Munar-Adrover etĀ al. (2011). However, based on the position of MYSO MSX G30 at the border of the confidence ellipse of the Fermi-LAT source, and the fact that the MYSO does not present any X-ray or centimeter radio emissions associated, we suggest that, it is very likely that the molecular outflows detected towards this MYSO are not energetic enough to be responsible for the -ray emission. Therefore, we conclude that it is very unlikely that the counterpart of the high-energy source is such a young protostar.
4.5 Is IRAS 184430231 generating the -ray emission of 4FGL J1846.90227?
@@
The source morphology at 3 and 5 GHz that we shown in Fig.ā10 together with the non-thermal nature of the radio continuum emission, which is in line with what was indicated by Irabor etĀ al. (2018), strongly suggest the presence of jets and accelerated particles in IRAS 184430231. Additionally, we discovered a red-shifted molecular outflow supporting the presence of such jets arising from the central source, and moreover, the X-ray morphology seems to encompass the outflow extension. The presence of jets and molecular outflows are in agreement with the detected H2O maser by Urquhart etĀ al. (2011).
@@
Taking into account that this source was previously catalogued as a PN candidate, our results indicate that IRAS 184430231 actually is a protoplanetary nebula (pPN), which is a short-lived transition (about yr) from the asymptotic giant branch (AGB) to the PN phase. There are not much cases that non-thermal radio continuum emission was observed towards this kind of objects (see the case of IRAS 15103-5754; SuĆ”rez etĀ al. 2015). On the other side, it is more common observing molecular outflows in pPNe (Cox etĀ al., 2000; Lorenzo etĀ al., 2021). The non-thermal radio emission in pPNe suggests that the emitting electrons arise at collisions between the fast and slow AGB winds that are observed predominantly on the front sides of the circumstellar shells (Bains etĀ al., 2009).
@@
Having discarded all the sources presented above as responsible of the -ray emission, taking into account the location of IRAS 184430231 (almost at the centre of the Fermi confidence ellipse), and mainly considering the presence of non-thermal jets accelerating particles, we wonder if IRAS 184430231 could be the counterpart of 4FGL J1846.90227.
@@
As mention by SuĆ”rez etĀ al. (2015) for IRAS 151035754, the radio spectral index between and measured in IRAS 18443-0231 is similar to those found in synchrotron emission in galaxies with active nuclei. Following the authors, in the energy conditions of PNe, mass-loss processes can produce collisionless, non-relativistic shocks, and electrons may undergo diffusive shock acceleration up to very high velocities (āfirst-order Fermi accelerationā). In such a scenario, in which particles are accelerated, and in addition the source seems to be related to some surrounding molecular molecular gas (see Sect.ā3.5), processes such as proton-proton collisions and relativistic Bremsstrahlung are likely to occur generating -rays.
Additionally, SuĆ”rez etĀ al. (2015) noted that the PN IRAS 15103-5754 presents non-thermal radio emission and a spectral index that has become flatter in the lapse of a few years. They attribute this flattening to the arising of a photoionized nebula. In our case, the radio spectral index obtained only from the THOR dataset towards IRAS 184430231 also could be revealing an arising of thermal emission. This would indicate that the very-high energy emission from pPNe should be a transient phenomenon.
In conclusion, we strongly suggest that IRAS 184430231 would be the responsible of the -ray emission. If this is the case, it would be the first reported pPN related to very high energy emission, and hence, multiwavelength dedicated observations, a follow up radio observations towards IRAS 184430231 and modelling are necessary to understand the mechanisms of -rays production in such a kind of source.
@@ 5 Summary and concluding remark
In previous works the nature of the Fermi-LAT source 4FGL J1846.90227 was discussed. In one of them it was suggested to be a blazar, and in other one it was mentioned the presence of a massive protostar within the Fermi confidence ellipse suggesting a possible association. Considering this discrepancy, in this work, we analysed several counterpart candidates to the -ray source based on a multiwavelength analysis.
@@
First of all, from JVLA data we determined the thermal nature of an uncatalogued extended radio source in the field, which indicates that it should be an HiiĀ region and not a supernova remnant, ruling out that such a source is responsible for the -ray emission. Using ALMA data, we confirmed that MSX6C G030.198100.1691 is a massive protostar with molecular outflow activity. However, since its peripheral location respect to the Fermi 95% confidence ellipse, and the lack of associated X-ray and centimeter radio continuum emissions, it is very unlikely that this massive protostar is responsible of the -ray emission.
@@
We found that sources SwXF4 J184651.6-022507 and SwXF4 J184650.7-022904, previously suggested to be blazars, do not meet the mid-IR colour-colour criteria for this kind of sources. Additionally, they lie almost at the galactic plane, which weakens the possibility of a blazar detection. Moreover, the second one is related to the 4XMM 184650.6022907 source, and based on XMM-Newton data, we found that this X-ray source, which lies closer to the centre of the Fermi 95% confidence ellipse, is clearly associated with the spectroscopic binary star BD-02 4739. The X-ray spectrum of 4XMM 184650.6022907 was well fitted by an optically thin thermal plasma model with a relatively low X-ray luminosity of about 1030 erg s-1. The soft thermal nature of the X-ray spectrum discards that 4XMM 184650.6022907 is associated with a blazar. In fact, the X-ray spectrum is consistent with a white-dwarf (WD) symbiotic system. Thus, taking into account the results of the analysis of the X-ray data and the optical/infrared SED, we conclude that BD02 4739 is a non-burning shell -type WD symbiotic system with a giant donor star. The only way that this source could be considered as the responsible for the -ray emission is that it had experimented a nova event. Despite of these new and interesting results about this source, the absence of an optical light curve indicating a nova event strongly suggests that BD-02 4739/4XMM 184650.6022907 is not associated with the high energy emission.
@@
Finally, we found that IRAS 18443-0231, catalogued as a PN candidate located at a distance of about 8 kpc, would be actually a protoplanetary nebula. IRAS 18443-0231 is the closest source in the region to the centre of the Fermi 95% confidence ellipse. We discovered extended and faint features related to the central peak of IRAS 184430231 at 3 and 5 GHz continuum emission and confirmed the non-thermal nature of the radio continuum (spectral index between and ). These findings indicate synchrotron emission very likely due to particles accelerated in jets. Additionally, we discovered a molecular outflow, supporting the presence of such jets, and we found that the source is surrounded by some molecular gas (traced by faint dust emission). IRAS 184430231 also presents hard X-ray extended emission, with the same inclination as the molecular outflow.
@@
Considering all the presented results, we conclude that the most promising source to be associated with 4FGL J1846.90227 is IRAS 184430231. The detection of synchrotron emission indicates that particles are accelerated at high velocities in the region, and considering the presence of some surrounding molecular gas, we suggest that processes such as proton-proton collisions and relativistic Bremsstrahlung are likely to occur. If this is the case, it would be the first reported protoplanetary nebula related to very high energy emission, and hence, multiwavelength dedicated observations and modelling are necessary to understand the mechanisms of -rays production in this kind of source.
@@ Acknowledgments
We thank the referee Dr. J. MartĆĀ for his comments and suggestions that helped us to improve this work. M.O., A.P., and S.P. are members of the Carrera del Investigador CientĆfico of CONICET, Argentina. This work was partially supported by the Argentina grants PIP 2021 11220200100012 and PICT 2021-GRF-TII-00061 awarded by CONICET and ANPCYT. This work is based on the following ALMA data: ADS/JAO.ALMA 2015.1.01312. ALMA is a partnership of ESO (representing its member states), NSF (USA) and NINS (Japan), together with NRC (Canada), MOST and ASIAA (Taiwan), and KASI (Republic of Korea), in cooperation with the Republic of Chile. The Joint ALMA Observatory is operated by ESO, AUI/NRAO and NAOJ. This publication makes use of VOSA, developed under the Spanish Virtual Observatory (https://svo.cab.inta-csic.es) project funded by MCIN/AEI/10.13039/501100011033/ through grant PID2020-112949GB-I00. VOSA has been partially updated by using funding from the European Unionās Horizon 2020 Research and Innovation Programme, under Grant Agreement nĀŗ 776403 (EXOPLANETS-A).
DATA AVAILABILITY
The data underlying this paper were accessed from:
- ā¢
- ā¢
- ā¢
- ā¢
- ā¢
- ā¢
- ā¢
- ā¢
- ā¢
The catalogues underlying this paper were accessed from:
- ā¢